a pilot of mass m can withstand a maximum apparent weight 6 times of mg.the minimum radius of curvature of vertical in which the aeroplane dives up from lowest point with a speed 504 km h is....solution

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
velocity at the top of the circle=v
v2=u2+2gh
=>v2=5042+4gr

At the top point the max upthrust=6mg
weight acting is mg
force acting downwards towards the center=mv2/r

Accordin to the FBD,
6mg-mg = mv2/r
=>5mg=(m/r)(5042+4gr)
=>r=5042/g
=>r=25401.6 m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)